# DRCT: Saving Image Super-Resolution away from Information Bottleneck

### v2

@@ 4.3 Same-task Progressive Training
In recent years, Progressive Training Strategy (PTS) [12, 54] have gained increased attention and can be seen as a method of fine-tuning. Compared to conventional training methods, PTS tends to converge model parameters to more desirable local-minima. Within the field of image processing, IPT [3] employs a variety of low-level tasks for pre-training, while EDT [22] utilizes different degradation levels within a single task. HAT [5] introduces the concept of same-task pre-training, where training on a larger and more diverse dataset like ImageNet [9] before fine-tuning on a specific dataset leads to improved super-resolution results. Lei et al. [49] proposed initially training a SISR network with L1-loss and then using L2-loss to eliminate artifacts, achieving better results on the PSNR metric. This concept has been widely adopted [58]. Our proposed Same-task Progressive Training (SPT), we first pre-trained DRCT on ImageNet to initialize model parameters and then fine-tuned on specific dataset with L1 loss as the optimization objective function,
$\text{Loss}_{1}=\left\|I_{HR}-I_{SR}\right\|_{1}$ (11)
@@
and finally use L2 loss to eliminate singular pixels and artifacts, therefore further archiving greater performance on PSNR metrics.
$\text{Loss}_{2}=\left\|I_{HR}-I_{SR}\right\|_{2}$ (12)
@@ 5 Experiment Results
DRCT (Ours) ×$\times 2$ DF2K 38.80 0.9646 35.02 0.9291 32.79 0.9071 34.58 0.9474 40.41 0.9816
@@ 5.1 Dataset
Our DRCT model is trained on DF2K, a substantial aggregated dataset that includes DIV2K [1] and Flickr2K [36]. DIV2K provides 800 images for training, while Flickr2K contributes a larger set of 2650 images. For the training input, we generate low-resolution versions of these images by applying a bicubic downsampling method with scaling factors of $2$, $3$, and $4$, respectively. To assess the effectiveness of our model, we conduct performance evaluations using well-known Image Super-Resolution benchmark datasets such as Set5 [2], Set14 [48], BSD100 [29], Urban100 [17], and Manga109 [30], which are standard in SISR tasks.
@@ 5.2 Implementation Details
The training process can be structured into three distinct phases, as ullustrated as Section 4-3. (1) pre-trained on ImageNet [9], (2) optimize the model on the given dataset, (3) L2-loss for PSNR enhancement. Throughout the training process, we use the Adam optimizer with $\beta_{1}$ = $0.9$, and $\beta_{2}$ = $0.999$ and train for $800k$ iterations in each stage. The learning rate is set to $42e-$ 4, the multi-step learning scheduler is also used. The learning rate is halved at the $300k$, $500k$, $650k$, $700k$, $750k$ iterations respectively. Weight decay is not applied and the batch-size is set to $32$. In the architecture of DRCT, the configuration of depth and width is maintained identically to that of SwinIR. To elaborate, both the number of RDFEG and SDRCB units are established at 6, and the channel number of intermediate feature maps is designated as 180. The attention head number and window size are set to 6 and 16 for window-based multi head self-attention (W-MSA). In terms of data preparation, HR patches with dimensions of $256$$\times$ $256$ pixels were extracted from the HR images. To improve the generalizability, we apply random horizontal flips and rotation augmentation.
@@ 5.3 Quantitative Results
For the evaluation, we use full RGB channels and ignore the $(2\times\text{scale})$ pixels from the border. PSNR and SSIM metrics are used to evaluation criteria. Table 1 presents the quantitative comparison of our approach and the state-of-the-art methods, including EDSR [25], RCAN [52], SAN [8], IGN [56], HAN [32], NLSN [31], SwinIR [24], CAT-A [6], DAT [7], as well as approaches using ImageNet pre-training, such as IPT [3], EDT [22] and HAT [5]. We can see that our method outperforms the other methods significantly on all benchmark datasets. In addition, the DRCT-L can bring further improvement and greatly expands the performance upper-bound of this task. Even with fewer model parameters and computational requirements, DRCT is also significantly greater than the state-of-the-art methods.
@@ 5.4 Visual Comparison
The visual comparisons displayed in Figure 5. For the selected images from the Urban100 [17], DRCT is effective in restoring structures, whereas other methods suffer from notably blurry effects. The visual results demonstrate the superiority of our approach.
Along with providing visualizations for the LAM [13], we compute the DI and GI. In scenarios where DRCT operates with fewer parameters (a topic for the next subsection), it achieves a higher DI. This outcome suggests that, after enhancing the receptive field through SDRCB, the model can leverage a greater amount of non-local information for SISR without the need for intricately designed W-MSA. Additionally, the GI of the proposed DRCT, being the lowest among the three, underscores its training process stability and its effectiveness in minimizing information loss.
There remains significant room for improvement in the way information loss is measured for SISR models. We conducted a detailed analysis of the G-index using ’image_092’ from Urban100, as shown in Figure 6. DRCT managed to achieve a lower score on the GI, indicating that the changes in the network’s feature maps are relatively stable. In contrast, The GI for HAT was significantly higher than the other two, which suggests that a lower GI value can effectively denote lesser information loss, whereas a higher GI might imply more drastic changes in feature maps. However, this doesn’t necessarily correlate directly with the effectiveness of measuring information loss.
Figure 6: The LAM [13] visualization. DRCT improves performance by enhancing the receptive field to mitigate the issue of spatial information loss in deeper layers of the network. (zoom in to better see the color-bar of feature maps.)
@@ 5.5 Model Complexity
To demonstrate the potential of our proposed method, we conducted further analysis on model complexity, with the results illustrated in Figure 7. From the figure, we can observe that the performance curves of the HAT and SwinIR models are approaching horizontal lines, suggesting that their model performance is nearing a bottleneck. Enhancing performance may require significantly more complex models or a greater number of parameters, potentially leading to prohibitive costs and practical implementation challenges. In contrast, our DRCT-L model outperforms HAT-L while saving 33% of model parameters. This indicates that stabilizing the training and inference processes within residual blocks through a dense-group is a viable strategy to prevent SISR models from encountering information bottlenecks.
@@ 5.6 NTIRE Image Super-Resolution (x4) Challenge
The dataset for the NTIRE 2024 Image Super-Resolution (x4) Challenge comprises three collections: DIV2K [1], Flickr2K [36], and LSDIR [23]. Specifically, the DIV2K dataset provides 800 pairs of high and low-resolution images for training. For validation, it offers 100 low-resolution images for the purpose of creating super-resolution counterparts, with the high-resolution versions to be made available at the challenge’s final stage. Additionally, the test dataset includes 100 varied images for generating their low-resolution equivalents. A self-ensemble strategy is used for testing-time augmentation (TTA) [35]. Our TTA methods include rotation, and horizontal and vertical flipping. We also conduct a model ensemble for fusing different reconstructed results by HAT [5] and the proposed DRCT. Our super-resolution (SR) model was entered into both the validation and testing phases of this Challenge, with the outcomes detailed in Table 2.
@@ 6 Conclusion
In this paper, we introduce the phenomenon of information bottlenecks observed in SISR models, where spatial information is lost as network depth increases during the forward-propagation process. This may lead to gradient vanishing and severe oscillations in the distribution of feature intensities, limiting the upper bound of model performance for the SISR task.
To address this issue, we present a novel Swin-transformer-based model named the Dense-residual-connected Transformer (DRCT). The design philosophy behind DRCT centers on stabilizing the forward-propagation process and enhancing the receptive field by incorporating dense connections within residual blocks. This approach reduces the loss of spatial information, addressing the information bottleneck issues that SISR models may encounter in deeper network layers.
As a result, the model can better focus on global spatial information and surpass existing state-of-the-art methods without the need for designing complex window attention mechanisms or increasing model parameters.
The extensive experiments demonstrate that DRCT surpasses previous methods in image super-resolution, indicating its effectiveness and the benefits of the proposed approach.

### v4

@@ 5.3 Same-task Progressive Training Strategy
In recent years, Progressive Training Strategy (PTS) [17, 25] has gained increased attention and can be seen as a method of fine-tuning. Compared to conventional training methods, PTS tends to converge model parameters to more desirable local minima. HAT [4] introduces the Same-task Pre-training, which aims to train the model on a large dataset like ImageNet [14] before fine-tuning it on a specific dataset, leading to improved SISR results. Lei et al. [57] proposed initially training a SISR network with L1-loss and then using L2-loss to eliminate artifacts, achieving better results on the PSNR metric. This has been widely adopted [64]. We proposed a Same-task Progressive Training Strategy (SPTS). At first, we pre-trained DRCT on ImageNet to initialize model parameters and then fine-tuned on specific datasets with L1 loss,
$\ell_{L1}=\left\|I_{HR}-I_{SR}\right\|_{1},$ (11)
@@
and finally use L2 loss to eliminate singular pixels and artifacts, therefore further archiving greater performance on PSNR metric.
$\ell_{L2}=\left\|I_{HR}-I_{SR}\right\|_{2}$ (12)
@@ 6 Experiment Results
DRCT (Ours) ×$\times 2$ DF2K 38.72 0.9646 34.96 0.9287 32.75 0.9071 34.54 0.9474 40.41 0.9814
@@ 6.1 Dataset
Our DRCT model is trained on DF2K, a substantial aggregated dataset that includes DIV2K [1] and Flickr2K [44]. DIV2K provides 800 images for training, while Flickr2K contributes 2650 images. For the training input, we generate LR versions of these images by applying a bicubic downsampling method with scaling factors of $2$, $3$, and $4$, respectively. To assess the effectiveness of our model, we conduct performance evaluations using well-known SISR benchmark datasets such as Set5 [2], Set14 [56], BSD100 [38], Urban100 [29], and Manga109 [24].
@@ 6.2 Implementation Details
The training process can be structured into three phases, as Section 4-3 illustrates. (1) pre-trained on ImageNet [14], (2) optimize the model on the given dataset, (3) L2-loss for PSNR enhancement. Throughout the training process, we use the Adam optimizer with $\beta_{1}$ = $0.9$, and $\beta_{2}$ = $0.999$ and train for $800k$ iterations in the first and second stages. The learning rate is set to $42e-$ 4, and the multi-step learning scheduler is also used. The learning rate is halved at the $300k$, $500k$, $650k$, $700k$, $750k$ iterations respectively. Weight decay is not applied, and the batch size is set to $32$. In the architecture of DRCT, the configuration of depth and width is maintained identically to that of HAT [4]. To elaborate, both the number of RDG and SDRCB units are established at 6, and the channel number of intermediate feature maps is designated as 180. The attention head number and window size are set to 6 and 16 for window-based multi-head self-attention (W-MSA). In terms of data preparation, HR patches with dimensions of $256$$\times$ $256$ pixels were extracted from the HR images. To improve the generalizability, we apply random horizontal flips and rotation augmentation.
@@ 6.3 Quantitative Results
For the evaluation, we use full RGB channels and ignore the $(2\times\text{scale})$ pixels from the border. PSNR and SSIM metrics are used to evaluation criteria. Table 1 presents the quantitative comparison of our approach and the state-of-the-art methods, including EDSR [35], RCAN [60], SAN [20], IGN [19], HAN [40], NLSN [39], SwinIR [34], CAT-A [5], DAT [6], as well as approaches using ImageNet pre-training, such as IPT [16], EDT [33] and HAT [4]. We can see that our method outperforms the other methods significantly on all benchmark datasets. In addition, the DRCT-L can bring further improvement and greatly expand the performance upper-bound on SISR tasks. Even with fewer model parameters and computational requirements, DRCT is also significantly greater than the state-of-the-art methods.
@@ 6.4 Visual Comparison
The visual comparisons displayed in Figure 4. For the selected images from Urban100 [29], DRCT is effective in restoring structures, whereas other methods suffer from notably blurry effects. The visual results demonstrate the superiority of our approach.
Along with providing visualizations for the LAM [27], we compute the Diffusion Index (DI), which is the attribution-based analysis. The DI reflects the range of involved pixels. A higher DI refers to a wider range of attention. In scenarios where DRCT used fewer parameters (which will be discussed in the next subsection), it achieves a higher DI. This outcome suggests that, after enhancing the receptive field through SDRCB, the model can leverage a long-range dependency and non-local information for SISR without the need for intricately designed W-MSA.
Figure 5: The LAM [27] visualization. DRCT improves performance by enhancing the receptive field to mitigate the issue of spatial information loss in deeper layers of the network.
@@ 6.5 Model Complexity
To demonstrate the potential of our proposed DRCT, we conducted further analysis on model complexity and performance.
Model efficiency. In Table 2, the proposed DRCT clearly requires fewer computational resources compared to HAT in terms of parameter size, multiply-add operations, memory requirements, and FLOPs. Specifically, when scaling up the model sizes of DRCT and HAT, DRCT-L surpasses HAT-L in all metrics.
Model performance. From Figure 6, we can observe that the performance curves of the HAT and SwinIR models are approaching horizontal lines, suggesting that the performance is nearing a bottleneck and its upper-bound, even if scaling up the model parameters.
This demonstrates that the design of DRCT, which incorporates dense-connections in the residual groups within a Swin-transformer-based model to stabilize the information flow, achieves convincing results with a reduced computational burden.
##\##Params. ##\##Multi-Adds. Forward or Backward pass FLOPs
HAT [4] 9.621M 11.22G 2053.42M 42.18G
DRCT 10.443M 5.92G 1857.55M 7.92G
HAT-L [4] 40.846M 76.69G 5165.39M 79.60G
DRCT-L 27.580M 9.20G 4278.19M 11.07G
Table 2: Model complexity analysis for (×4) SISR on Urban100.
@@ 6.6 NTIRE Image Super-Resolution (x4) Challenge
The dataset for the NTIRE 2024 Image Super-Resolution (x4) Challenge [7] comprises three collections: DIV2K [1], Flickr2K [44], and LSDIR [23]. Specifically, the DIV2K dataset provides 800 pairs of HR and LR images for training. The LR images are obtained from the HR images after bicubic downsampling with specific scaling factor. For validation, it offers 100 LR images for the purpose of creating SR images, with the HR versions to be made available at the challenge’s final stage. Additionally, the test dataset includes 100 varied LR images. The self-ensemble strategy is used for testing-time augmentation (TTA) [43]. Our TTA methods include random rotation, and horizontal and vertical flipping. We also conducted a model ensemble strategy for fusing different reconstructed results by HAT [4] and the proposed DRCT to eliminate the annoying artifacts and improve final SR quality. Our SISR model was entered into both the validation and testing phases of this challenge, with the detailed in Table 3.
@@ 7 Conclusion
In this paper, we introduce the phenomenon of information bottlenecks observed in SISR models, where spatial information is lost as network depth increases during forward propagation. This may lead to information loss when limiting the upper bound of model performance for the SISR task, which requires detailed spatial information and context-aware processing.
To address these issues, we present a novel Swin-transformer-based model, Dense-residual-connected Transformer (DRCT). The design philosophy behind DRCT centers on stabilizing the information flow and enhancing the receptive fields by incorporating dense-connections within residual blocks, combining the shift-window attention mechanism to adaptively capture global information.
As a result, the model can better focus on global spatial information and surpass existing state-of-the-art methods without the need for designing sophisticated window attention mechanisms or increasing model parameters. The experiment results have demonstrated the efficacy of the proposed DRCT, indicating its effectiveness and the potential for future work related to SISR tasks.